### v1

# On the limitation of evaluating machine unlearning using only a single training seed

Abstract
Machine unlearning (MU) aims to remove the influence of certain data points from a trained model without costly retraining. Most practical MU algorithms are only approximate and their performance can only be assessed empirically. Care must therefore be taken to make empirical comparisons as representative as possible. A common practice is to run the MU algorithm multiple times independently starting from the same trained model. In this work, we demonstrate that this practice can give highly non-representative results because – even for the same architecture and same dataset – some MU methods can be highly sensitive to the choice of random number seed used for model training. We therefore recommend that empirical comparisons of MU algorithms should also reflect the variability across different model training seeds.
3040506070SSDLFSSDUNSIRBadTeacherRandomLabelsUnlearning methodRetain-set accuracy [%]One training seed(common practice)Multiple training seeds(our recommendation)
(a) CIFAR100 – pine class.
255075SSDLFSSDUNSIRBadTeacherRandomLabelsUnlearning methodForget-set accuracy [%]
(b) CIFAR20 – sea (sub) class.
Figure 1: Retain-set accuracy (Panel 1(a)) and forget-set accuracy (Panel 1(b)), when using one training seed (combined with eleven unlearning seeds) versus eleven training seeds (each paired with one unlearning seed). Black lines represent the 25, 50 and 75 % quantiles of the accuracy of Retrain (the gold standard) across eleven seeds; whiskers extend to the full data range for clarity. This figure shows that the performance of some unlearning methods (e.g., SSD and LFSSD) is highly sensitive to choice of the training seed. Thus, empirical evaluation of unlearning methods should be based on multiple training seeds.
1 Introduction
Machine unlearning. The targeted removal of the influence of certain training data on the parameters of machine-learning models is often needed, e.g., for regulatory compliance (e.g., GDPR [10]), elimination of biased or corrupted samples, or mitigation of data poisoning attacks. These concerns extend to information retrieval systems [7, 8, 13]. However, the exact (‘gold-standard’) solution to the machine-unlearning (MU) problem, which involves retraining the model from scratch without the to-be-forgotten data (called Retrain), is typically infeasible. Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, e.g., Random-Labels [5], UNSIR [12], Bad-Teacher [2], SSD [3] or LFSSD [4].
Evaluating unlearning methods. MU algorithms are often difficult to analyse mathematically. Their evaluation must then be done empirically. This calls for empirical experiments be designed carefully so as to provide confidence that their findings generalise beyond the specific setting of the study (e.g., architecture or dataset). For instance, common practice is to run the unlearning algorithm multiple times independently – each using a different random-number generator (RNG) seed, henceforth called an unlearning seed. One then assesses the empirical distribution (across unlearning seeds) of unlearning metrics such as accuracy on the data to be forgotten (forget-set accuracy) or accuracy on the remaining data (retain-set accuracy).
Contributions. While comparisons across multiple unlearning seeds are common, most works only consider a single training seed, i.e., they take the same single trained model as the starting point of unlearning. However, as we point out in this work, this practice can give highly non-representative results (see Figure 1). That is, even for the same architecture and dataset, some MU methods can be highly sensitive to the choice of training seed. This issue seems to particularly affect the deterministic MU methods: SSD and LFSSD. For a fixed trained model, such deterministic methods always return exactly the same unlearned model – and hence exactly the same values of any unlearning metric – no matter what unlearning seed is used. However, Figure 1 illustrates that unlearning metrics for SSD and LFSSD are actually highly variable across different training seeds. Specifically, our contributions are as follows:
  1. 1.
    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for SSD and LFSSD.
  2. 2.
    We explain the counter-intuitive fact that an unlearning metric can exhibit a smaller variability across multiple training seeds than for a single, fixed training seed.
  3. 3.
    We recommend evaluating unlearning methods across multiple training seeds.
2 Methods
Generic unlearning evaluation. Algorithm 1 outlines a generic empirical evaluation of some MU method for a given architecture and given dataset.
Algorithm 1 Generic empirical evaluation of unlearning methods
Number of distinct training seeds: $I\geq 1$.
Number of distinct unlearning seeds per training seed: $J\geq 1$.
Distinct training seeds $s_{1},\dotsc,s_{I}$.
Distinct unlearning seeds $\smash{\tilde{s}}_{i,1},\dotsc,\smash{\tilde{s}}_{i,J}$, for each $1\leq i\leq I$.
for $i=1,\dotsc,I$ do
  $M_{i}\leftarrow\operatorname{train}(\text{seed}=s_{i})$ $\triangleright$ trained model
  for $j=1,\dotsc,J$ do
   $U_{i,j}\leftarrow\operatorname{unlearn}(M_{i},\text{seed}=\smash{\tilde{s}}_{i,j})$ $\triangleright$ unlearned model
  end for
end for
Unlearning performance across $\{U_{i,j}\mid 1\leq i\leq I,1\leq j\leq J\}$.
Here, the model is trained II times independently using a different training seed; each of the trained models is then a starting point for $J$ (conditionally) independent runs of the unlearning approaches. Finally, the unlearning performance is evaluated by calculating some unlearning metric (e.g., forget-set accuracy or retain-set accuracy) across all $I\times J$ unlearned models.
Two special cases. This work discusses the following special cases:
  1. A
    Common practice is to evaluate unlearning across
    • only $I=1$ training seed;
    • $J\geq 1$ unlearning seeds (e.g., J=3J=3[5], $J=5$ [2], $J=10$ [4, 3]).
  2. B
    Our recommendation is to evaluate unlearning across
    • $I>1$ training seeds (ideally, II should be much larger than 11);
    • $J\geq 1$ unlearning seeds per training seed.
Variance-decomposition analysis. One might intuitively expect that A underestimates the variability of the chosen unlearning metric. However, this is not necessarily true as we now show. In fact – depending on the single training seed – it is possible that A overestimates variability.
To see this, let $S$ and $\smash{\tilde{S}}$ be randomly chosen training and unlearning seeds, respectively. Further, for some given architecture, unlearning method, and retain and forget sets, let Z=h(S,S~)Z=h($\smash{\tilde{S}}$) be the value of some metric for assessing unlearning under a model first trained using training seed $S$ and then unlearned using unlearning seed $\smash{\tilde{S}}$ (e.g., $Z$ could be the accuracy on the forget set or retain set). In the following, we only focus on the variance of the unlearning metric, but the issue affects the entire distribution of the unlearning metric and thus also applies to other moments, e.g., the mean.
  • Common practice (A) leads to an approximation of the conditional distribution of $Z$ given the single training-seed value $s_{1}$, i.e., it yields estimates of the following conditional variance:
[…]
  • Our recommendation (B) leads to an approximation of the unconditional distribution of $Z$, i.e., it yields estimates of the following (unconditional) variance, where the second identity is a standard variance decomposition:
[…]
3 Experiments
In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extends to other classes and datasets.
As datasets, we chose CIFAR-100 full-class [9], CIFAR-20 full-class, and CIFAR-20 sub-class [2]. For CIFAR-20 sub-class, the 100 classes in CIFAR-100 were grouped into 20 superclasses. We used a ResNet-18 [6] architecture, randomly initialised and trained from scratch on each dataset. Experiments were run on an NVIDIA A100-SMX4-80GB GPU. Code for reproducing the experiments – based on https://github.com/if-loops/selective-synaptic-dampening with minimal modifications – is available at https://github.com/jtlan90/evaluating-machine-unlearning-using-only-a-single-training-seed/.
We compare the empirical distributions of the unlearning metrics: forget-set accuracy and retain-set accuracy, in terms of 2-Wasserstein distance. For each metric, these empirical distributions are based on the $I\times J=11$ realisations obtained from A (Algorithm 1 with $I=1$ and $J=11$) and from B (Algorithm 1 with $I=11$ and $J=1$). Figure 2 shows the results. Therein, each boxplot is based on the 2-Wasserstein distances for each of the unlearning classes.
Retain-set accuracyForget-set accuracyCIFAR100(full class)CIFAR20(full class)CIFAR20(sub class)SSDLFSSDUNSIRBadTeacherRandomLabelsSSDLFSSDUNSIRBadTeacherRandomLabels0510150102030051015Unlearning method2-Wasserstein distance($I=1$ vs $I=11$)
Figure 2: 2-Wasserstein distance of the empirical distribution (per unlearning class) of the unlearning metrics: retain-set accuracy and forget-set accuracy, under the common-practice approach ($I=1$ & $J=11$) vs our recommendation ($I=11$ & $J=1$). This figure illustrates that the results from Figure 1 extend to other unlearning classes and datasets, i.e., that considering only a single training seed can lead to a highly non-representative unlearning-metric distribution, especially for SSD and LFSSD.
4 Conclusion
Summary. We have demonstrated that the performance of machine unlearning (MU) methods – as measured by the accuracy on the retain and forget sets – can be highly sensitive to the initial model weights as determined by training seed. Thus, using only a single training seed for evaluating MU methods (which is common practice in the literature) can yield misleading conclusions about their efficacy or robustness. Results for another MU metric: Membership inference attack (MIA) [1, 11], were qualitatively similar and were thus omitted in the interest of brevity.
Recommendations. Based on our findings, MU evaluations should use multiple training seeds (>1>1) to ensure results are robust.
Limitations and extensions. Our work illustrates the potential pitfall of evaluating machine unlearning using only a single training seed. However, we note that for each dataset and model, our experiments compare only one instance of the ‘common-practice’ MU evaluation approach (i.e., a single training seed but multiple unlearning seeds) with only one instance of our recommended evaluation approach (i.e., multiple training seeds). A more comprehensive analysis (which would require multiple instances of each approach) is beyond the scope of this work. We leave this to future work.

### v2

@@
 Unlearning performance across $\{U_{i,j}\mid 1\leq i\leq I,1\leq j\leq J\}$.
-Here, the model is trained II times independently using a different training seed; each of the trained models is then a starting point for $J$ (conditionally) independent runs of the unlearning approaches.
+Here, the model is trained $I$ times independently using a different training seed; each of the trained models is then a starting point for $J$ (conditionally) independent runs of the unlearning approaches.
 Finally, the unlearning performance is evaluated by calculating some unlearning metric (e.g., forget-set accuracy or retain-set accuracy) across all $I\times J$ unlearned models.
@@
     Our recommendation is to evaluate unlearning across
-    • $I>1$ training seeds (ideally, II should be much larger than 11);
+    • $I>1$ training seeds (ideally, $I$ should be much larger than $1$);
     • $J\geq 1$ unlearning seeds per training seed.

### v3

@@
 In this work, we demonstrate that this practice can give highly non-representative results because – even for the same architecture and same dataset – some MU methods can be highly sensitive to the choice of random number seed used for model training.
+We illustrate that this is particularly relevant for MU methods that are deterministic, i.e., which always produce the same result when started from the same trained model.
 We therefore recommend that empirical comparisons of MU algorithms should also reflect the variability across different model training seeds.
@@
 Black lines represent the 25, 50 and 75 % quantiles of the accuracy of Retrain (the gold standard) across eleven seeds; whiskers extend to the full data range for clarity.
-This figure shows that the performance of some unlearning methods (e.g., SSD and LFSSD) is highly sensitive to choice of the training seed.
-Thus, empirical evaluation of unlearning methods should be based on multiple training seeds.
+This figure shows that the performance of the deterministic unlearning methods: SSD and LFSSD, is highly sensitive to choice of the training seed.
+Thus, empirical evaluation of (especially such deterministic) unlearning methods should be based on multiple training seeds.
 1 Introduction
 Machine unlearning.
-The targeted removal of the influence of certain training data on the parameters of machine-learning models is often needed, e.g., for regulatory compliance (e.g., GDPR [10]), elimination of biased or corrupted samples, or mitigation of data poisoning attacks.
-These concerns extend to information retrieval systems [7, 8, 13].
+The targeted removal of the influence of certain training data on the parameters of machine-learning models [Machine_Unlearning_Cao] is often needed, e.g., for regulatory compliance (e.g., GDPR [mantelero2013eu]), elimination of biased or corrupted samples, or mitigation of data poisoning attacks.
+These concerns extend to information retrieval systems [hou2024neural, hou2024neural_corrective, zhang2022machine].
 However, the exact (‘gold-standard’) solution to the machine-unlearning (MU) problem, which involves retraining the model from scratch without the to-be-forgotten data (called Retrain), is typically infeasible.
-Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, e.g., Random-Labels [5], UNSIR [12], Bad-Teacher [2], SSD [3] or LFSSD [4].
+Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, e.g., Random Labels [GolatkarEternalsunshineoftheSpotless202], UNSIR [Tarun_2024_UNSIR], Bad Teacher [chundawat2023badteachinginduceforgetting], SSD [Foster_Schoepf_Brintrup_2024] or LFSSD [foster2024loss].
 Evaluating unlearning methods.
@@
 Their evaluation must then be done empirically.
-This calls for empirical experiments be designed carefully so as to provide confidence that their findings generalise beyond the specific setting of the study (e.g., architecture or dataset).
-For instance, common practice is to run the unlearning algorithm multiple times independently – each using a different random-number generator (RNG) seed, henceforth called an unlearning seed.
+This calls for empirical experiments to be designed carefully so as to provide confidence that their findings generalise beyond the specific setting of the study (e.g., architecture or dataset).
+Common practice [chundawat2023badteachinginduceforgetting, Chundawat_2023, graves2021amnesiac, GolatkarEternalsunshineoftheSpotless202, Tarun_2024_UNSIR, Foster_Schoepf_Brintrup_2024, foster2024loss, ForgettingOutsidetheboxgolatkar, mixedprivforgettinggolatkar, bourtoule2021machine, unrollingsgdanvith, deepregressionunlearningtarun23a, lee2025unlearningcomparatorvisualanalytics] is to run stochastic unlearning algorithms (e.g., UNSIR, Random Labels or Bad Teacher) multiple times independently – each using a different random-number generator (RNG) seed, henceforth called an unlearning seed, to account for their inherent variability.
 One then assesses the empirical distribution (across unlearning seeds) of unlearning metrics such as accuracy on the data to be forgotten (forget-set accuracy) or accuracy on the remaining data (retain-set accuracy).
+In contrast, deterministic MU algorithms, such as SSD and LFSSD, are only run once [Foster_Schoepf_Brintrup_2024, foster2024loss] because they always return exactly the same unlearned model – and hence exactly the same values of any unlearning metric – no matter what unlearning seed is used.
 Contributions.
-While comparisons across multiple unlearning seeds are common, most works only consider a single training seed, i.e., they take the same single trained model as the starting point of unlearning.
-However, as we point out in this work, this practice can give highly non-representative results (see Figure 1).
-That is, even for the same architecture and dataset, some MU methods can be highly sensitive to the choice of training seed.
-This issue seems to particularly affect the deterministic MU methods: SSD and LFSSD.
-For a fixed trained model, such deterministic methods always return exactly the same unlearned model – and hence exactly the same values of any unlearning metric – no matter what unlearning seed is used.
-However, Figure 1 illustrates that unlearning metrics for SSD and LFSSD are actually highly variable across different training seeds.
+While comparisons across multiple unlearning seeds are common, many works use a single training seed, i.e., a single trained model as the starting point of unlearning.
+In this work, we demonstrate that such a strategy can lead to highly non-representative results because – even for a given architecture and data set – MU methods can be highly sensitive to the choice of training seed.
+Our work thus confirms similar conclusions recently drawn in [triantafillou2024makingprogressunlearningfindings, cadet2025deep].
+Notably, our work differs from [cadet2025deep, triantafillou2024makingprogressunlearningfindings] by demonstrating that this issue particularly affects the deterministic MU methods: SSD and LFSSD.
+As illustrated in Figure 1, unlearning metrics for SSD and LFSSD – although constant across unlearning seeds for a given training seed – exhibit disproportionately high variability across different training seeds.
 Specifically, our contributions are as follows:
   1. 1.
-    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for SSD and LFSSD.
+    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for the deterministic methods: SSD and LFSSD.
   2. 2.
-    We explain the counter-intuitive fact that an unlearning metric can exhibit a smaller variability across multiple training seeds than for a single, fixed training seed.
+    In particular, we show that using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
   3. 3.
-    We recommend evaluating unlearning methods across multiple training seeds.
+    We explain the counter-intuitive fact that an unlearning metric can exhibit a smaller variability across multiple training seeds than for a single, fixed training seed (when the latter is combined with multiple unlearning seeds).
 2 Methods
@@
     • only $I=1$ training seed;
-    • $J\geq 1$ unlearning seeds (e.g., J=3J=3[5], $J=5$ [2], $J=10$ [4, 3]).
+    • $J\geq 1$ unlearning seeds (e.g., J=3J=3[GolatkarEternalsunshineoftheSpotless202], $J=5$ [chundawat2023badteachinginduceforgetting], $J=10$ [foster2024loss, Foster_Schoepf_Brintrup_2024]).
   2. B
-    Our recommendation is to evaluate unlearning across
+    Our recommendation reflecting that of [cadet2025deep] is to evaluate unlearning across
     • $I>1$ training seeds (ideally, $I$ should be much larger than $1$);
@@
 3 Experiments
-In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extends to other classes and datasets.
-As datasets, we chose CIFAR-100 full-class [9], CIFAR-20 full-class, and CIFAR-20 sub-class [2].
+In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extend to other classes and datasets.
+As datasets, we chose CIFAR-100 full-class [krizhevsky2009learning], CIFAR-20 full-class, and CIFAR-20 sub-class [chundawat2023badteachinginduceforgetting].
 For CIFAR-20 sub-class, the 100 classes in CIFAR-100 were grouped into 20 superclasses.
-We used a ResNet-18 [6] architecture, randomly initialised and trained from scratch on each dataset.
+We used a ResNet-18 [resnet] architecture, randomly initialised and trained from scratch on each dataset.
 Experiments were run on an NVIDIA A100-SMX4-80GB GPU.
-Code for reproducing the experiments – based on https://github.com/if-loops/selective-synaptic-dampening with minimal modifications – is available at https://github.com/jtlan90/evaluating-machine-unlearning-using-only-a-single-training-seed/.
+Code for reproducing the experiments – based on https://github.com/if-loops/selective-synaptic-dampening with minimal modifications – is available (anonymised for review) at https://anonymous.4open.science/r/evaluating-machine-unlearning-using-only-a-single-training-seed-88CC.
 We compare the empirical distributions of the unlearning metrics: forget-set accuracy and retain-set accuracy, in terms of 2-Wasserstein distance.
@@
 Summary.
-We have demonstrated that the performance of machine unlearning (MU) methods – as measured by the accuracy on the retain and forget sets – can be highly sensitive to the initial model weights as determined by training seed.
-Thus, using only a single training seed for evaluating MU methods (which is common practice in the literature) can yield misleading conclusions about their efficacy or robustness.
-Results for another MU metric: Membership inference attack (MIA) [1, 11], were qualitatively similar and were thus omitted in the interest of brevity.
-Recommendations. Based on our findings, MU evaluations should use multiple training seeds (>1>1) to ensure results are robust.
+We have confirmed a recent finding from [cadet2025deep] that the performance of machine unlearning (MU) methods can be highly sensitive to the initial model weights as determined by training seed.
+Thus, using only a single training seed for evaluating MU methods (a common practice in the literature) can yield misleading conclusions about their efficacy or robustness.
+Complementary to [cadet2025deep], our work shows that: (a) this issue can be of particular concern when evaluating MU methods that are deterministic; (b) using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
+Empirical results shown in this work are for the MU metrics: accuracy on the retain and forget sets; results for another MU metric: Membership inference attack (MIA) [MIAunlearningprivacychenetal, MIAshokrietal], were qualitatively similar and were thus omitted in the interest of brevity.
+Recommendations. Based on our work we echo the recommendations of [cadet2025deep] that MU evaluations should use multiple training seeds (>1>1) to ensure results are robust.
 Limitations and extensions.
-Our work illustrates the potential pitfall of evaluating machine unlearning using only a single training seed.
+Our work specifically illustrates the potential pitfall of evaluating machine unlearning using only a single training seed.
 However, we note that for each dataset and model, our experiments compare only one instance of the ‘common-practice’ MU evaluation approach (i.e., a single training seed but multiple unlearning seeds) with only one instance of our recommended evaluation approach (i.e., multiple training seeds).

### v4

@@
 We therefore recommend that empirical comparisons of MU algorithms should also reflect the variability across different model training seeds.
-3040506070SSDLFSSDUNSIRBadTeacherRandomLabelsUnlearning methodRetain-set accuracy [%]One training seed(common practice)Multiple training seeds(our recommendation)
-(a) CIFAR100 – pine class.
-255075SSDLFSSDUNSIRBadTeacherRandomLabelsUnlearning methodForget-set accuracy [%]
-(b) CIFAR20 – sea (sub) class.
-Figure 1: Retain-set accuracy (Panel 1(a)) and forget-set accuracy (Panel 1(b)), when using one training seed (combined with eleven unlearning seeds) versus eleven training seeds (each paired with one unlearning seed).
-Black lines represent the 25, 50 and 75 % quantiles of the accuracy of Retrain (the gold standard) across eleven seeds; whiskers extend to the full data range for clarity.
-This figure shows that the performance of the deterministic unlearning methods: SSD and LFSSD, is highly sensitive to choice of the training seed.
-Thus, empirical evaluation of (especially such deterministic) unlearning methods should be based on multiple training seeds.
 1 Introduction
@@
 However, the exact (‘gold-standard’) solution to the machine-unlearning (MU) problem, which involves retraining the model from scratch without the to-be-forgotten data (called Retrain), is typically infeasible.
-Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, e.g., Random Labels [GolatkarEternalsunshineoftheSpotless202], UNSIR [Tarun_2024_UNSIR], Bad Teacher [chundawat2023badteachinginduceforgetting], SSD [Foster_Schoepf_Brintrup_2024] or LFSSD [foster2024loss].
+Instead, a number of approximate MU methods have been developed to enable models to forget specific data whilst maintaining overall performance, i.e., Random Labels [GolatkarEternalsunshineoftheSpotless202], UNSIR [Tarun_2024_UNSIR], Bad Teacher [chundawat2023badteachinginduceforgetting], SSD [Foster_Schoepf_Brintrup_2024] or LFSSD [foster2024loss] in image classification and Gradient Ascent [halimi2023federatedunlearningefficientlyeraseGA], FedRemove [FedRemoveYuan2023] and FedEraser [FedEraser2021Liu] in Federated Online Learning to Rank (FOLTR) unlearning.
 Evaluating unlearning methods.
@@
 In this work, we demonstrate that such a strategy can lead to highly non-representative results because – even for a given architecture and data set – MU methods can be highly sensitive to the choice of training seed.
-Our work thus confirms similar conclusions recently drawn in [triantafillou2024makingprogressunlearningfindings, cadet2025deep].
-Notably, our work differs from [cadet2025deep, triantafillou2024makingprogressunlearningfindings] by demonstrating that this issue particularly affects the deterministic MU methods: SSD and LFSSD.
+Our work thus confirms similar conclusions recently drawn in [triantafillou2024makingprogressunlearningfindings] and [cadet2025deep].
+Notably, our work differs from [cadet2025deep] and [triantafillou2024makingprogressunlearningfindings] by demonstrating that this issue particularly affects the deterministic MU methods: SSD and LFSSD as well as exploring this issue within federated online learning to rank.
 As illustrated in Figure 1, unlearning metrics for SSD and LFSSD – although constant across unlearning seeds for a given training seed – exhibit disproportionately high variability across different training seeds.
@@
   1. 1.
-    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for the deterministic methods: SSD and LFSSD.
+    We illustrate that single-training-seed based evaluation of unlearning methods produces non-representative results, especially for the deterministic methods SSD and LFSSD.
   2. 2.
-    In particular, we show that using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
+    We demonstrate that using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
   3. 3.
-    We explain the counter-intuitive fact that an unlearning metric can exhibit a smaller variability across multiple training seeds than for a single, fixed training seed (when the latter is combined with multiple unlearning seeds).
+    We explain the counter-intuitive fact that an unlearning metric can exhibit smaller variability across multiple training seeds than for a single, fixed training seed (when the latter is combined with multiple unlearning seeds).
+  4. 4.
+    We explore how this issue affects task domains beyond image classification [cadet2025deep], to those such as federated online learning to rank (FOLTR).
+3040506070SSDLFSSDUNSIRBadTeacherRandomLabelsUnlearning methodRetain-set accuracy [%]One training seed(common practice)Multiple training seeds(our recommendation)
+(a) CIFAR100 – pine class.
+255075SSDLFSSDUNSIRBadTeacherRandomLabelsUnlearning methodForget-set accuracy [%]
+(b) CIFAR20 – sea (sub) class.
+Figure 1: Retain-set accuracy (Panel 1(a)) and forget-set accuracy (Panel 1(b)), when using one training seed (combined with eleven unlearning seeds) versus eleven training seeds (each paired with one unlearning seed).
+Black lines represent the 25, 50 and 75% quantiles of the accuracy of Retrain (the gold standard) across eleven seeds; whiskers extend to the full data range for clarity.
+This figure shows that the performance of the deterministic unlearning methods: SSD and LFSSD, is highly sensitive to choice of the training seed.
+Thus, empirical evaluation of (especially such deterministic) unlearning methods should be based on multiple training seeds.
 2 Methods
-Generic unlearning evaluation. Algorithm 1 outlines a generic empirical evaluation of some MU method for a given architecture and given dataset.
+Generic unlearning evaluation. Algorithm 1 outlines a generic empirical evaluation of some MU method for a given architecture and dataset.
 Algorithm 1 Generic empirical evaluation of unlearning methods
@@
     • only $I=1$ training seed;
-    • $J\geq 1$ unlearning seeds (e.g., J=3J=3[GolatkarEternalsunshineoftheSpotless202], $J=5$ [chundawat2023badteachinginduceforgetting], $J=10$ [foster2024loss, Foster_Schoepf_Brintrup_2024]).
+    • $J\geq 1$ unlearning seeds (e.g., $J=3$ [GolatkarEternalsunshineoftheSpotless202], $J=5$ [chundawat2023badteachinginduceforgetting], $J=10$ [foster2024loss, Foster_Schoepf_Brintrup_2024]).
   2. B
@@
 3 Experiments
-In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extend to other classes and datasets.
+In this section, we demonstrate that the differences between the common-practice evaluation approach (A) and our recommended approach (B) shown in Figure 1 extends to other classes, datasets and tasks.
+All experiments were ran on a shared lab computer on utilising a single NVIDIA A100-SMX4-80GB GPU.
+Retain-set accuracyForget-set accuracyCIFAR100(full class)CIFAR20(full class)CIFAR20(sub class)SSDLFSSDUNSIRBadTeacherRandomLabelsSSDLFSSDUNSIRBadTeacherRandomLabels0510150102030051015Unlearning method2-Wasserstein distance($I=1$ vs $I=11$)
+Figure 2: 2-Wasserstein distance of the empirical distribution (per unlearning class) of the unlearning metrics: retain-set accuracy and forget-set accuracy, under the common-practice approach ($I=1$ & $J=11$) vs our recommendation ($I=11$ & $J=1$).
+This figure illustrates that the results from Figure 1 extend to other unlearning classes and datasets, i.e., that considering only a single training seed can lead to a highly non-representative unlearning-metric distribution, especially for SSD and LFSSD.
+3.1 Image Classification Task
 As datasets, we chose CIFAR-100 full-class [krizhevsky2009learning], CIFAR-20 full-class, and CIFAR-20 sub-class [chundawat2023badteachinginduceforgetting].
@@
 We used a ResNet-18 [resnet] architecture, randomly initialised and trained from scratch on each dataset.
-Experiments were run on an NVIDIA A100-SMX4-80GB GPU.
-Code for reproducing the experiments – based on https://github.com/if-loops/selective-synaptic-dampening with minimal modifications – is available (anonymised for review) at https://anonymous.4open.science/r/evaluating-machine-unlearning-using-only-a-single-training-seed-88CC.
+The following unlearning methods were examined within this context – Random Labels [GolatkarEternalsunshineoftheSpotless202], UNSIR [Tarun_2024_UNSIR], Bad Teacher [chundawat2023badteachinginduceforgetting], SSD [Foster_Schoepf_Brintrup_2024] or LFSSD [foster2024loss].
 We compare the empirical distributions of the unlearning metrics: forget-set accuracy and retain-set accuracy, in terms of 2-Wasserstein distance.
@@
 Therein, each boxplot is based on the 2-Wasserstein distances for each of the unlearning classes.
-Retain-set accuracyForget-set accuracyCIFAR100(full class)CIFAR20(full class)CIFAR20(sub class)SSDLFSSDUNSIRBadTeacherRandomLabelsSSDLFSSDUNSIRBadTeacherRandomLabels0510150102030051015Unlearning method2-Wasserstein distance($I=1$ vs $I=11$)
-Figure 2: 2-Wasserstein distance of the empirical distribution (per unlearning class) of the unlearning metrics: retain-set accuracy and forget-set accuracy, under the common-practice approach ($I=1$ & $J=11$) vs our recommendation ($I=11$ & $J=1$).
-This figure illustrates that the results from Figure 1 extend to other unlearning classes and datasets, i.e., that considering only a single training seed can lead to a highly non-representative unlearning-metric distribution, especially for SSD and LFSSD.
+3.2 Federated Online Learning to Rank (FOLTR) Task
+In FOLTR multiple clients collectively train a ranking model to rank documents for search queries while learning from simulated user interactions (clicks).
+Unlearning aims to remove the influence where malicious clients or incorrect data poison the federated ranking model.
+The following unlearning methods were examined – finetune [finetuneFOLTRZhang2023], Gradient Ascent [halimi2023federatedunlearningefficientlyeraseGA], FedRemove [FedRemoveYuan2023] and FedEraser [FedEraser2021Liu].
+We compare common-practice ($I=1$ & $J=10$) (Figure 3(a)) vs our recommendation ($I=10$ & $J=1$) (Figure 3(b)).
+We utilise the MQ2007 [LETORdataset2013] dataset.
+A linear ranker [linearranker2021foltr], randomly initialised and trained from scratch in a federated online learning setting with 10 clients on the dataset.
+Three click behaviour models (Perfect, Navigational, and Informational) were employed to simulate different user interaction patterns, following [foltrrepTao].
+Evaluation is performed as demonstrated in Figure 3 using normalised discounted cumulative gain at position 10 (NDCG@10), which measures ranking quality by comparing the system’s ranking to the ideal ranking based on document relevance labels.
+NDCG@10 is normalised to [0, 1], where higher values indicate better unlearning performance.
+(a) One Training Seed
+(b) Multiple Training Seeds
+Figure 3: FOLTR on MQ2007 model poison unlearning scenario.
+Offline NDCG@10 (evaluated on held-out test set after each training/unlearning epoch), scores are plotted over 2000 epochs.
+The first 1000 epochs represent the training phase, epochs 1000-2000 represent the unlearning phase.
+Thin transparent lines show individual, runs thick solid lines represent the mean trajectory and shaded regions represent standard deviation.
 4 Conclusion
@@
 Thus, using only a single training seed for evaluating MU methods (a common practice in the literature) can yield misleading conclusions about their efficacy or robustness.
-Complementary to [cadet2025deep], our work shows that: (a) this issue can be of particular concern when evaluating MU methods that are deterministic; (b) using multiple unlearning seeds cannot compensate for the lack of multiple training seeds.
-Empirical results shown in this work are for the MU metrics: accuracy on the retain and forget sets; results for another MU metric: Membership inference attack (MIA) [MIAunlearningprivacychenetal, MIAshokrietal], were qualitatively similar and were thus omitted in the interest of brevity.
+Complementary to [cadet2025deep], our work shows that: (a) this issue can be of particular concern when evaluating MU methods that are deterministic; (b) using multiple unlearning seeds cannot compensate for the lack of multiple training seeds; (c) demonstrated this issue extends beyond image classification MU methods.
+Empirical results shown in this work are for the MU metrics: accuracy on the retain and forget sets (Iamge) and NDCG@10 (FOLTR); results for another MU metric: Membership inference attack (MIA) [MIAunlearningprivacychenetal, MIAshokrietal], were qualitatively similar and were thus omitted in the interest of brevity.
 Recommendations. Based on our work we echo the recommendations of [cadet2025deep] that MU evaluations should use multiple training seeds (>1>1) to ensure results are robust.
-Limitations and extensions.
+5 Limitations
 Our work specifically illustrates the potential pitfall of evaluating machine unlearning using only a single training seed.
@@
 A more comprehensive analysis (which would require multiple instances of each approach) is beyond the scope of this work.
-We leave this to future work.
+Future work could include additional tasks such as unlearning in large language models, and additional evaluations against more datasets within the FOLTR task; however this additional experimentation would require large computational resources.
+Appendix A Data and Code availability
+Code for reproducing both experiments – based on https://github.com/if-loops/selective-synaptic-dampening (Image) and https://github.com/Iris1026/Unlearning-for-FOLTR (FOLTR) with minimal modifications – is available (anonymised for review) at https://anonymous.4open.science/r/evaluating-machine-unlearning-using-only-a-single-training-seed-88CC.